Consider the economy of Hicksonia

a) The consumption function is given by: C=300+0.6(Y-T). The investment function is: I=700-80r. Government purchases and taxes are both 500. For this economy, graph the IS curve for r changing from 0 to 8

b) The money demand function in Hicksonia is (M/P)^d=Y-200r. The money supply M is 3000 and the price level P is 3. For this economy, graph the LM curve for r changing from 0 to 8.

c) Find the equilibrium interest rate r and the equilibrium level of income Y.

d) Suppose that government purchases are increased from 500 to 700. How does the IS curve shift? What are the new equilibrium interest rate and level of income?

e) Suppose instead that the money supply is raised from 3000 to 4500. How does the LM cruve shift? What are the new eqquilibrium interest rate and level of income?

f) With the initial level values for monetary and fiscal policy, suppose that the price level rises from 3 to 5. What hpapnes? What are the new equilibrium interest rate and level of income?

g) Derive and graph and equation for the aggregate demand curve.What happens to this aggregate demand curve if fiscal or monetary policy changes, as in part (d) and (e)?
Answer #1

(a)

IS curve equation is:

Y = C + I + G

Y = 300 + 0.6(Y - 500) + 700 - 80r + 500 [Since G = T = 500]

= 1500 + 0.6Y - 300 - 80r

(1 - 0.6)Y = 1200 - 80r

0.4Y = 1200 - 80r

Y = 3000 - 200r [Dividing both sides by 0.4] ..... (1) [Equation of IS]

IS
r Y
0 3000
1 2800
2 2600
3 2400
4 2200
5 2000
6 1800
7 1600
8 1400

(b) LM Equation is:

(M/P)d = M

(3000 / 3) = Y - 200r

1000 = Y - 200r

Y = 1000 + 200r ....... (2) [LM Equation]

LM
r Y
0 1000
1 1200
2 1400
3 1600
4 1800
5 2000
6 2200
7 2400
8 2600

(c)

Equilibrium is obtained when IS = LM

3000 - 200r = 1000 + 200r

400r = 2000

r = 5

Y = 1000 + 200r = 1000 + (200 x 5) = 1000 + 1000 = 2000

(d)

If G increases from 500 to 700 (by 200), from part (a) we get

0.4Y = 1200 - 80r + 200 = 1400 - 80r

Y = 3500 - 200r ....... (3) [Revised IS curve]

Equating (3) with LM (2):

3500 - 200r = 1000 + 200r

400r = 2500

r = 6.25

Y = 1000 + (200 x 6.25) = 1000 + 1250 = 2250

NOTE: First 4 sub-questions are answered in full.